OpenStudy (anonymous):

The equation of the line in point-slope form is y-6=3/7(x+8) Write in general form using the formula Ax+By=C

OpenStudy (jlastino):

y-6=(3/7)x+8 Multiply both sides by 7 7(y-6)= 7((3/7)x+8) 7y-42=3x+56 3x-7y+98=0

OpenStudy (anonymous):

its not 98

OpenStudy (jlastino):

Oh yeah my bad it should be in general from 3x-7y=-98

OpenStudy (anonymous):

no its 66 but i need to kno why it stays a positive when u put 3x on the other side of the equal sign